# Mamo: a Mathematical Modeling Benchmark with Solvers

### v2

@@ Abstract
Mathematical modeling involves representing real-world phenomena, systems, or problems using mathematical expressions and equations to analyze, understand, and predict their behavior. Given that this process typically requires experienced experts, there is an interest in exploring whether Large Language Models (LLMs) can undertake mathematical modeling to potentially decrease human labor. To evaluate of LLMs in mathematical modeling, we introduce a new benchmark, Mamo, that transcends traditional result-oriented assessments. Unlike conventional methods that primarily assess LLMs based on the accuracy of solutions to mathematical problems, our approach offers deeper insight into the modeling process itself. By focusing on the processes LLMs undertake rather than the correctness of their final solutions, Mamo pioneers a novel evaluation paradigm. This shift underscores the importance of understanding the inherent modeling capabilities of LLMs, paving the way for a more nuanced and comprehensive analysis of their problem-solving strategies. Our work marks a significant advancement in the field, suggesting a new direction for future research by emphasizing the evaluation of LLMs’ modeling processes over the mere correctness of answers. This benchmark not only facilitates a better understanding of LLMs’ mathematical modeling capabilities but also sets a new standard for evaluating their performance in complex problem-solving scenarios.
@@ 1 Introduction
Recent advancements in Large Language Models have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks Ouyang et al. (2022); OpenAI (2023); Nijkamp et al. (2022); Tang et al. (2024). However, the domain of mathematics remains a critical aspect of LLM evaluation Wei et al. (2022). This focus not only gauges LLMs’ proficiency in comprehending and addressing mathematical challenges but also serves as a profound indicator of their underlying abilities in abstract conceptualization and logical reasoning, among other intellectual faculties.
@@
Natural language captures human complexity and nuance, while mathematical language excels in precision and modeling the natural world (Giordano et al., 2013). Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise. However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
In this endeavor, solvers [^1] In optimization, solvers are algorithms used to find the best solution to a problem. They search through possible options to maximize or minimize an objective under given constraints if any. By using solvers, The challenge becomes problem formulation under a given solver or a solver set. becomes critically important as it standardizes the modeling output and could provide a way to falsifiably validate the produced answers. Recent studies have already combined LLMs with solvers Yang et al. (2023); Feng et al. (2023); Pan et al. (2023). In the field of optimization, AhmadiTeshnizi et al. (2023b) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities.
We thus develope Mamo, a new benchmark that leverages solvers to assess the mathematical modeling prowess of LLMs. Our approach shifts from conventional outcome-focused evaluation to process-focused evaluation, offering in-depth insights into the LLMs’ problem-solving strategies. We concentrate on the models’ mathematical modeling skills, delegating the task of solving abstract problems to solvers to circumvent errors arising from the models’ computational capabilities.
@@
The contributions of this paper are as below: (1) We introduce a new benchmarking strategy to assess mathematical modeling through Solvers, facilitating a rigorous evaluation process. (2) We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions (With a total of 1209 meticulously curated questions), including ordinary differential equations and optimization problems within linear programming and mixed-integer linear programming frameworks.
2 Related Work
Recent studies have made significant strides in applying LLMs to mathematical problem-solving, introducing innovative datasets for comprehensive evaluation. Frieder et al. (2024) and Yuan et al. (2023) have focused on datasets that test LLMs across various math problems and arithmetic expressions, respectively. Lewkowycz et al. (2022) explored training models using natural language paired with LaTeX-formatted mathematical content from arXiv, enhancing syntactic understanding. Zhang et al. (2024) developed the CARP dataset for computation-intensive challenges, while He et al. (2024) introduced OlympiadBench, a bilingual and multimodal benchmark for competition-level mathematical reasoning and proofs. Liu et al. (2024) constructed MathBench, which is designed with a carefully structured difficulty hierarchy and focuses on assessing the understanding of theoretical knowledge in LLMs. AhmadiTeshnizi et al. (2023a) constructed NLP4LP, a benchmark comprising 52 linear programming (LP) and mixed-integer linear programming (MILP) problems. These efforts collectively push the boundaries of LLMs in mathematical cognition, showcasing the growing complexity and specificity of tasks LLMs are being trained to tackle.
The pursuit of enhancing and devising novel solvers has seen a variety of research efforts, with a notable focus on the integration of LLMs. Yang et al. (2023) investigated the potential of employing LLMs as direct solvers, pioneering the exploration of LLMs’ autonomous problem-solving capabilities. He-Yueya et al. (2023) introduced a hybrid approach that marries an LLM with an external symbolic solver, specifically for equation solving, marking a significant advancement in combining neural and symbolic computation. Furthermore, Pan et al. (2023) developed LOGIC-LM, a cutting-edge framework that fuses LLMs with symbolic solvers, aimed at enhancing the solving of logical problems. In a closely related study, AhmadiTeshnizi et al. (2023a) introduced OptiMUS, an LLM-based agent explicitly crafted for formulating and solving optimization problems, thus demonstrating the broad applicability and potential of LLMs in tackling intricate computational challenges. These endeavors underscore the innovative directions being explored in the realm of solver development, leveraging the robust capabilities of LLMs.
3 Background
3.1 Mathematical Modeling
Giordano et al. (2013) emphasize the role of mathematical models as intermediaries that transform real-world problems into structured mathematical forms, facilitating the discovery of solutions. Despite LLMs’ proven competencies across various tasks, their capacity for developing mathematical models remains uncertain. This involves two key phases: model formulation, which demands a thorough comprehension of the problem, and model resolution, typically addressed by computational solvers. To accurately gauge LLMs’ potential in this domain, our strategy focuses on their ability to construct mathematical models, entrusting calculations to specialized solvers. This approach highlights LLMs’ creativity and generalization skills, simultaneously capitalizing on solvers’ computational efficiency.
Motivationsto Benchmark Mathematical Modeling for LLMs   We evaluate mathematical models for two reasons. The first reason is out of curiosity: We aim to investigate whether LLMs can do what expert humans can. Mathematical modeling is considered a highly advanced capability that we believe experts possess. This is a beneficial test for exploring the boundaries of LLMs’ capabilities and to what extent they can approach human intelligence. On the other hand, in terms of evaluating the models themselves, most current benchmarks, especially those related to mathematics like GSM8K Cobbe et al. (2021), etc., show that the performance has become saturated. Even a 7B model Yu et al. (2023b); Liu et al. (2023) (sometimes with a verifier-like model Yu et al. (2023a)) can achieve a accuracy exceeding 80%, closely approaching that of GPT 4. Therefore, we believe there is a need to construct more challenging mathematical datasets. Traditional mathematical modeling is a labor-intensive, custom process reliant on human expertise, raising the question: Can large language models automate or assist in this process?
3.2 Using Solvers in Mathematical Modeling
The model generated by the LLM is then subjected to verification against real-world data to assess its validity, ensuring it addresses the problem logically and sensibly. The solver, an algorithmic tool designed to find solutions to mathematical models, is also involved in this process and it is also beneficial for verification.
Different types of solvers are specialized for various classes of problems, such as optimization solvers. However, they share the common property: They are all algorithmic tools designed to find solutions to mathematical models.
Rationale of Using solver in Mathematical Modeling for LLMs These solvers operationalize the abstract constructs created by LLMs, translating them into tangible outputs that can be analyzed and validated against real-world data and phenomena. The solver is a crucial instrument that resolves the model presented by the LLM, and by comparing the solution it provides with the correct answer, we assess the accuracy of the mathematical model generated by the LLM.
4 Philosophy of Mamo Benchmark
4.1 Methodology: exact answer verification with solvers
We utilize solver to resolve this mathematical model. The solver operates at the final state of the question-solving process, meaning it directly computes the final answer from the model without intermediate steps. By executing the solver and juxtaposing its output ($\widehat{A}$) with the correct answer (A) to the original problem, we can verify the accuracy of the Language Model’s mathematical model(${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$), since typically the wrong ${\widehat{M}}$ will result in wrong ($\widehat{A}$), like the examples in Appendix  G. This final-state approach ensures that our assessment is squarely focused on the Language Model’s capacity to model the problem, abstracting from any additional computational steps that might otherwise obscure the evaluation of the model’s validity.
@@ 4.1 Methodology: exact answer verification with solvers
4.2 Scope of Mathematical Modeling
We construct the category of Mathematical Modeling by referring to Giordano et al. (2013). In the context of Natural Language we propose a framework in mathematical modelling in LLM, specified in solving natural language problem, see Table 7 in Appendix F.
@@ 4.2 Scope of Mathematical Modeling
Given the broad spectrum of mathematical modeling and the availability of specific solvers, our benchmark is meticulously designed to encompass domains where these tools can be effectively utilized. This section outlines the rationale for our focus on certain areas of mathematical modeling, based on solver availability and the clarity of the modeling process.
@@
Availability of Sophisticated Solvers: Advanced optimization solvers, such as COPT Ge et al. (2023), along with Python libraries for solving ordinary differential equations (ODEs), offer a strong foundation for testing the capabilities of Large Language Models (LLMs). These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
Challenges in Modeling Changes and Differences, Proportionality and Geometric Similarity: Although simpler tools like calculators could address problems related to modeling changes and differences, proportionality and geometric similarity, the lack of a defined, structured modeling process complicates the implementation of our final-state-approach benchmark. Our objective is to assess the LLMs’ skill in developing comprehensive mathematical models that adhere to a solvable framework. The absence of a distinct modeling procedure in these areas impedes a standardized and objective assessment of an LLM’s modeling efficiency.
@@
Following the considerations mentioned above, we have developed our benchmark in the fields of optimization, and ODE, due to the lack of advanced and universal Partial Differential Equations (PDE) solvers. In the optimization segment, our focus is on Linear Programming (LP) and Mixed-Integer Linear Programming (MILP) problems. Compared to NLP4LP, introduced by AhmadiTeshnizi et al. (2023b) and focusing exclusively on optimization, our benchmark includes a larger number of problems. This benchmark set is highly scalable, with potential future expansions into nonlinear optimization and probabilistic modeling. Additionally, as new and advanced solvers emerge, it could be expanded into areas like Partial Differential Equations (PDE) and beyond.
5 Data Collection
5.1 Data Selection and Source Credibility
The dataset for our benchmark consists of a blend of manually selected and GPT-synthetized questions. All items within the data set have been subjected to a meticulous review process conducted by qualified collectors and domain experts to ensure their validity and relevance. The data collection and synthesis process is conducted by a team of qualified individuals with extensive expertise in various mathematical disciplines  [^2] See the details of the annotator qualifications in Appendix C.
Manually Selected Questions
The manually selected questions are high quality and reliable. (1)Source and Reliability: When doing manually selecting, we refer to the textbooks, exercises and the solutions in ODE and optimization area(Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce & DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Giordano et al., 2014; Bertsimas & Tsitsiklis, 1997; Hurlbert, 2010). The selecting process focus on deriving the mathematical model in the questions and assign a reasonable scenario. The source ensure the correctness of the answer of each question. (2) Transformation: Each question is meticulously transformed to fit the benchmark’s criteria(the properties in Section 5.3), which may involve assigning new scenarios or adapting the question and its answer to align with our testing methodology.
@@ 5.1 Data Selection and Source Credibility
The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems. This approach not only tests the model’s ability to handle mathematical reasoning but also its capacity to contextualize mathematical principles within real-world situations. See the example in Appendix K.
5.2 Data Quality Checking
In the ODE part, we employed GPT-4 to confirm the solvability of newly generated questions, filtering out invalid ones. A thorough evaluation identified questions lacking information, with incorrect answers, and unclear statements. Consequently, problematic questions were deleted or corrected, resulting in a refined dataset. In the optimization section, particularly with the easy LP part, invalid questions were filtered out, and those with incorrect answers or misleading descriptions were corrected. In the complex LP part, questions formulated with wrong mathematical models were corrected. The summary of the review process is presented in Table 1.
Category Initial Questions Filtered Final Questions
ODE Part 383 37 346
Easy LP Part 688 48 640
Complex LP Part 211 0 211
@@ 5.2 Data Quality Checking
Reviewer Accuracy Rate
Reviewer 1 90.0%
Reviewer 2 84.0%
Reviewer 3 88.0%
Reviewer 4 74.0%
Table 2: Accuracy of reviewers.
For the cross-review process, 50 questions from the ODE dataset were selected for deeper analysis by four independent reviewers. Reviewers evaluated the questions, marking responses as either numeric answers or "error" if the question was problematic. The effectiveness of this review was quantified using metrics like Cohen’s Kappa to measure inter-rater reliability and accuracy rates for individual reviewers. The average Cohen’s Kappa was 0.60, indicating moderate to substantial agreement among reviewers, with individual Kappa values ranging from 0.50 to 0.71. Accuracy rates are moderately high as shown in Tables 2. These results highlight both the effectiveness of the review process and areas where alignment and training can improve consistency and accuracy among reviewers. See the review details in Appendix B.
5.3 Guidance of Data Collection
The collection of problems for our benchmark are governed by specific criteria designed to evaluate the mathematical modeling capacity of Large Language Models (LLMs). The guidelines ensure that problems are appropriate for our final-state approach and yield numerical answers suitable for automated verification. The following are the key considerations in our benchmark guidance:
@@ 5.3 Guidance of Data Collection
Final-State Approach Solvability  Problems should be structured so that they can be solved by a solver in a final-state approach, meaning the solution process should not require intermediate steps once the mathematical model has been formulated by the LLM.
@@
Unified and Numerical Answers: To handle potentially various valid solutions for one problem: For Ordinary Differential Equations (ODEs), questions should focus on determining the value of the original function at a specific time $t_{0}$. For optimization problems, questions should aim to find the optimal value. These constraints facilitate the testing accuracy and the convenience of verifying potentially multiple valid solutions to a single problem. Besides, the answers to all questions must be numerical, allowing for clear-cut evaluation and comparison with the solver’s output.
@@
Significant Figures and Precision  Each question will explicitly state the required significant figures or the level of numerical precision for the answer. This clarity ensures that the LLM’s output can be properly evaluated against the solver’s results, maintaining consistency and rigor in the assessment of the model’s accuracy.
@@
Real-World Problem Context  Questions should be framed as real-world problems without explicitly stating the underlying mathematical model, challenging the LLM to demonstrate its capacity for abstraction and application in practical scenarios.

### v3

# LLMs for Mathematical Modeling: Towards Bridging the Gap between Natural and Mathematical Languages

@@ Abstract
Large Language Models (LLMs) have demonstrated strong performance across various natural language processing tasks, yet their proficiency in mathematical reasoning remains a key challenge. Addressing the gap between natural and mathematical language requires advanced reasoning capabilities, approaching those of Artificial General Intelligence (AGI). However, the evaluation remains challenging, as perfectly representing reality is inherently elusive, and traditional methods like manual or direct comparison of mathematical statements (Ramamonjison et al., 2023) are insufficient for assessing true modeling ability. We propose a process-oriented framework to evaluate LLMs’ ability to construct mathematical models, using solvers to compare outputs with ground truth. Introducing Mamo, a benchmark with 1,209 questions covering ordinary differential equations, linear programming, and mixed-integer linear programming, we enable automatic evaluation of modeling accuracy. The results show that existing LLMs struggle with complex mathematical modeling tasks, with larger models demonstrating superior performance, while open-source models remain competitive in simpler cases but still fall short of proprietary models in more challenging problems.
LLMs for Mathematical Modeling: Towards Bridging the Gap between Natural and Mathematical Languages
Xuhan Huang1, Qingning Shen1, Yan Hu1thanks: Corresponding author: huyan@cuhk.edu.cn, Anningzhe Gao2, Benyou Wang1 1 The Chinese University of Hong Kong, Shenzhen 2 Shenzhen Research Institute of Big Data xuhanhuang,qingningshen@link.cuhk.edu.cn, huyan@cuhk.edu.cn anningzhegao@gmail.com, wangbenyou@cuhk.edu.cn
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have garnered widespread interest, demonstrating remarkable capabilities across a broad spectrum of natural language processing tasks (Ouyang et al., 2022; OpenAI, 2023; Nijkamp et al., 2022; Tang et al., 2024). However, the domain of mathematical reasoning remains a critical aspect of the LLM evaluation (Wei et al., 2022). This focus gauges LLMs’ proficiency in solving mathematical challenges and reveals their underlying abilities in abstract reasoning and logic.
@@
Natural language reflects the complexity and subtlety of human communication, while mathematical language provides accuracy and is ideal for representing the natural world (Giordano et al., 2013). Bridging their gap requires high-level cognitive skills akin to Artificial General Intelligence (AGI), a task at which LLMs show promise. Recent studies have already combined LLMs with solvers (Feng et al., 2023; Pan et al., 2023). In the field of optimization, AhmadiTeshnizi et al. (2023) introduced OptiMUS, an LLM-based agent that uses a solver to address optimization problems. The key task for LLM in OptiMUS involves first generating a mathematical formulation of a real-world problem, followed by writing Python code to invoke a solver. This implies the potential for testing LLM’s modeling capabilities. However, the challenge lies in the evaluation of modeling, as perfectly representing reality may be inherently elusive.
To address this, we propose a new benchmarking strategy that evaluates mathematical modeling by utilizing solvers and answer verification. The core philosophy is to use solvers to resolve mathematical models and assess their accuracy by comparing the solver’s output with the ground truth. Additionally, we introduce Mamo, a novel benchmark designed to assess the mathematical modeling capabilities of LLMs, leveraging solvers to shift the focus from conventional outcome-based evaluations to a more process-oriented approach. This strategy is particularly rational because it allows us to isolate and scrutinize the entire modeling process, from formulation to solution, providing a more holistic and rigorous evaluation of an LLM’s mathematical reasoning abilities.
The contributions of this paper are as follows:
  • We introduce an automatic evaluation framework to assess mathematical modeling through solvers, enabling exact answer matching.
@@
  • We have developed a new benchmark, named Mamo, specifically tailored for the evaluation of mathematical modeling capabilities. Mamo encompasses a wide array of modeling questions (with a total of 1,209 meticulously curated questions), including ordinary differential equations (ODEs) and optimization problems within linear programming and mixed-integer linear programming frameworks.
2 An Evaluation Framework for Mathematical Modeling
2.1 Definition: Mathematical Modeling
Mathematical modeling is widely considered an advanced skill typically possessed by experts. According to Giordano et al. (2013), mathematical models act as intermediaries that convert real-world problems into structured mathematical forms, facilitating the discovery of solutions. As illustrated in Figure 1, we aim to map a natural language question $Q$ into a mathematical model $M$, which is expressed in mathematical language as follows:
$f:Q\rightarrow M.$
By solving this model, we can obtain an answer $A$ to the original problem $Q$. Traditional mathematical modeling is a labor-intensive, highly specialized process that depends heavily on human expertise.
2.2 Gap between Natural and Mathematical Languages
Natural language captures the complexity and nuance of human thought, while mathematical language excels in precision and in modeling natural phenomena (Giordano et al., 2013). A single statement in natural language may correspond to different mathematical representations, and conversely, distinct natural language statements may map to the same mathematical formulation.
For example, consider the natural language statement: The car starts with an initial velocity of zero. This can correspond to at least two different mathematical representations:
Mathematical formulation 1: $y^{\prime}(0)=0$
Mathematical formulation 2: $v(0)=0$.
Conversely, these two mathematical expressions may also correspond to natural language 2: The population growth rate is initially zero.
Bridging the gap between natural language and mathematical language requires advanced cognitive abilities, a task where LLMs have shown significant potential. However, evaluating the effectiveness of mathematical modeling remains challenging, as achieving a perfect representation of reality through models is often unattainable.
Figure 1: $Q$ is the natural language problem, while $M_{1}$,$M_{2}$ are different mathematical models.
Challenges in Evaluation
Due to the complex mapping between natural language and mathematical language, a real-world problem $Q$ may correspond to multiple equivalent mathematical models, such as $M_{1},M_{2},M_{3},\dots$. As illustrated in Figure 1, for a given problem $Q$, we have two distinct mathematical models, $M_{1}$ and $M_{2}$, both of which correctly describe the same real-world problem and are considered equivalent in the context of $Q$. Traditionally, the evaluation of mathematical models, including determining their equivalence, has been performed by human experts, a process that is both labor-intensive and time-consuming. In optimization modeling area, Ramamonjison et al. (2023) directly compare mathematical statements by standardizing variable names and evaluating accuracy based on declaration matches. While this approach addresses differences in variable naming, the abstraction of variables from natural language context, which is critical in mathematical modeling, was not examined.
@@ 2.2 Gap between Natural and Mathematical Languages
2.3 Evaluating Modeling Using Solvers and Answer Matching
The Evaluation Framework
To facilitate the automatic evaluation of mathematical models, we propose an evaluation framework for mathematical modeling based on exact answer verification using solvers, as illustrated in Figure 2. A solver is an algorithmic tools designed to find solutions to various classes of problems. For instance, Gurobi is a well-known optimization solver (Gurobi Optimization, LLC, 2024), while Python packages (Virtanen et al., 2020) and MATLAB (Inc., 2022) are commonly used for solving ODEs. Solvers are essential for operationalizing abstract models, denoted as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$, generated by large language models (LLMs), translating them into concrete solutions $\widehat{A}$ that can be analyzed and validated against real-world data. For a given real-world problem and its corresponding answer pair $(Q,A)$, LLMs generate a mathematical model, denoted ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$. The solver resolves this model, producing a solution $\widehat{A}$, which is then compared to the correct answer $A$. This framework directly evaluates the LLM’s problem-solving capability, with any inaccuracies in $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$ typically leading to incorrect solutions $\widehat{A}$. By focusing on the final output, we avoid additional computational steps that could obscure the evaluation of the LLM’s performance.
The Philosophy
The philosophy behind our evaluation framework is goal-driven, based on the idea that the ultimate aim of mathematical modeling is to produce accurate solutions to real-world problems. We prioritize the correctness of the solution $\widehat{A}$ over the intermediate steps of model formulation. By comparing the solver’s output with the correct answer, we evaluate the effectiveness of the LLM’s generated model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$. The assumption here is that if the model is well-constructed, it will lead to an accurate solution.
Despite its goal-driven structure, the framework retains a process-oriented emphasis on the modeling capability of LLMs. Unlike traditional benchmarks that treat answer correctness as an isolated metric, our approach explicitly evaluates the intermediate step of model formulation. The evaluative focus lies on whether the LLM constructs a mathematical model $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}$ that faithfully translates the real-world problem into mathematical language. Thus, while the solver’s answer validates the outcome, the core objective remains to assess the modeling process.
This dual perspective allows us to maintain a scalable yet principled evaluation framework. While the solver’s output remains the validation mechanism, the primary objective is to assess the LLM’s ability to conceptualize and formulate mathematical problems rather than simply solving predefined mathematical tasks (more discussions are provided in Appendix M).
3 Mamo Benchmark
To implement this framework, we developed a new benchmark, Mamo, specifically for mathematical modeling.
3.1 Scope of Mamo Benchmark
@@
Given the broad spectrum of mathematical modeling Giordano et al. (2013) (see Appendix G) and the availability of specific solvers, our benchmark is meticulously designed to encompass domains of ODEs and linear programming problems. This section outlines the rationale behind our focus on these areas, driven by the criteria of solver availability and result verifiability.
Availability of Sophisticated Solvers:
@@
Advanced optimization solvers, such as COPT (Ge et al., 2023), along with Python libraries for solving ODEs, offer a strong foundation for testing the capabilities of LLMs. These tools enable a detailed and automated evaluation of the LLMs’ abilities to abstract, formulate, and accurately solve mathematical problems within these specific domains.
The Straightforward Verifiability of Results
While calculators address basic problems in these areas, the absence of a structured modeling approach hinders the assessment of LLMs’ ability to create solvable mathematical models.
@@
Building on these considerations, we developed a benchmark focused on optimization and ODE, given the lack of universal PDE solvers. Compared to NLP4LP, introduced by AhmadiTeshnizi et al. (2023) and focusing exclusively on optimization, our benchmark includes a larger number of problems, with potential expansions into nonlinear optimization, probabilistic modeling, and PDE as new solvers emerge.
3.2 Data Selection
Our benchmark dataset comprises manually selected and GPT-generated questions, curated by a team with advanced mathematical education, including Ph.D. holders, and an average GPA of 3.89/4.0, ensuring strong analytical expertise. Details on the collectors’ qualifications are provided in Appendix C.
Source Reliability
For manual selection, we referenced textbooks and solutions in ODE and optimization (Thomas et al., 2016; Stewart et al., 2014; Lial et al., 2017; Braun, 1993; Boyce and DiPrima, 2012; Giordano et al., 2014; Zill, 2013; Bertsimas and Tsitsiklis, 1997; Hurlbert, 2010). We focused on deriving mathematical models and assigning realistic scenarios, ensuring answer accuracy. Each question was adapted to meet the benchmark criteria (Section 3.3) by adjusting scenarios or solutions as needed.
@@ 3.2 Data Selection
Data synthesis starts by creating mathematical constructs with random parameters, followed by answer computation to establish ground truth. GPT-4 then generates real-life scenarios, translating these models into context-rich problems, testing both mathematical reasoning and real-world contextualization. See the example in Appendix N.
3.3 Guidance of Data Collection
Figure 3: $Q$ is the natural language problem (may be ODE problem or optimization problem), where y is a solution function in ODE problems.
Category Initial Questions Filtered Corrected Final Questions
ODE 383 37 41 346
Easy LP 688 36   8 652
Complex LP 211   0 40 211
@@ 3.3 Guidance of Data Collection
Our benchmark dataset follows specific criteria to ensure the problems are suitable for automated verification. Figure 3 provides examples of these criteria, with further explanations presented in Appendix T.
@@
Final-State Approach Solvability  Problems should be structured so that a solver can directly compute the solution once the LLM formulates the mathematical model. For example, asking for the optimal value of an optimization problem is solvable using a ’final-state approach’, whereas transforming an ODE solution requires extra steps and is not solvable this way.
@@
Unified and Numerical Answers
For ODEs, questions should seek the function’s value at a specific time $t_{0}$, and for optimization problems, the optimal value, ensuring a unified and numerical answer for clear comparison with the solver’s output.
@@
Significant Figures and Precision
Each question will explicitly state the required significant figures or the level of numerical precision for the answer.
@@
Real-World Problem Context Questions should be framed as real-world problems without explicitly stating the underlying mathematical model.
Reviewer Reviewer 1 Reviewer 2 Reviewer 3 Reviewer 4
Reviewer 1 - 0.67 0.71 0.56
Reviewer 2 0.67 - 0.65 0.50
Reviewer 3 0.71 0.65 - 0.52
Reviewer 4 0.56 0.50 0.52 -
Table 2: Inter-reviewer Cohen’s Kappa scores.
Reviewer Accuracy Rate (%)
Reviewer 1 90.0
Reviewer 2 84.0
Reviewer 3 88.0
Reviewer 4 74.0
Table 3: Accuracy of reviewers.
3.4 Data Quality Checking
In the ODE section, we utilized GPT-4 to verify the solvability of generated questions as stated in Property 1, filtering out any invalid questions. A detailed evaluation identified issues such as missing information, incorrect answers, and unclear phrasing, leading to the removal or correction of problematic questions and resulting in a refined dataset. In the optimization section, specifically in the Easy_LP part, invalid questions were eliminated, and those with incorrect answers or misleading descriptions were corrected. For the Complex_LP section, questions involving incorrect mathematical models were revised (see an example in Appendix J). A summary of the review process is provided in Table 1.
For the cross-review process, 50 questions from the ODE dataset were selected for in-depth analysis by four independent reviewers  [^1] The hourly wage is approximately 20 US dollars, with total compensation of around 473 US dollars.. Reviewers were chosen based on their qualifications, including an average GPA of at least 3.5/4.0 in fundamental math courses, completion or enrollment in the Ordinary Differential Equation course, and meeting English proficiency standards.
Reviewers evaluated the questions, classifying responses as either numerical answers or "error" if the question was problematic. The effectiveness of this review was assessed using Cohen’s Kappa to measure inter-rater reliability, along with individual accuracy rates. As shown in Table 3, the pairwise Cohen’s Kappa scores had an average of 0.60, reflecting moderate to substantial agreement among reviewers, with individual scores ranging from 0.50 to 0.71. Accuracy rates were moderately high, as detailed in Tables 3. For additional details on the review, see Appendix B.